Beryllium(II) forms a complex with acetylacetone (166.2 g/mol). Calculate the molar absorptivity of the complex, given that a 2.25 ppm solution has a transmittance of 37.5% when measured in a 1.00-cm cell at 295 nm, the wavelength of maximum absorption
1
Expert's answer
2020-05-19T08:58:23-0400

Solution.

"2.25 \ ppm = 2.25 \times 10^{-3} \ g/L"

"Cm = \frac{m}{M \times V}"

"Cm = \frac{2.25 \times 10^{-3}}{166.2} = 1.35 \times 10^{-5} M"

"A = \xi \times l \times Cm"

"\xi = \frac{A}{l \times Cm}"

"A = lg[\frac{I0}{I}]"

"A = lg[\frac{100}{37.5}] = 0.426"

"\xi = \frac{0.426}{1 \times 1.35 \times 10^{-5}} = 3.16 \times 10^4 \ \frac{L}{mol \times cm}"

Answer:

"\xi = 3.16 \times 10^4 \ \frac{L}{mol \times cm}"